OpenStudy (mathmath333):

10 students can do a job in 8 days, but on the starting day , two of them informed that they are not coming. by what fraction will the number of days required for the doing the whole work gets increased ? a.) 4/5 b.) 3/8 c.) 3/4 d.) 1/4

ganeshie8 (ganeshie8):

`10 students can do a job in 8 days` so 1 student can do the same job in 80 days that means 8 students can do it in 10 days

ganeshie8 (ganeshie8):

it takes 2 extra days to finish the job : \[\dfrac{2}{8}\]

OpenStudy (mathmath333):

how u came over the \(\dfrac{2}{8}\)

ganeshie8 (ganeshie8):

it asks for a fraction of "number of days increased"

ganeshie8 (ganeshie8):

the reference is the original 8 days, so..

OpenStudy (mathmath333):

ok \(\dfrac{\rightarrow 2}{8}\) how does the \(2\) comes

ganeshie8 (ganeshie8):

as per original plan, it has to take `8` days to finish the job but now it is going to take `10` days that just means 2 extra days

OpenStudy (mathmath333):

ok thnx